# refrences.bib

### v1

@@
Abstract
This paper, in the first step, develops the system of bipolar fuzzy relational equations (FRE) to the most general case where the bipolar FREs are defined by an arbitrary continuous t-norm. Also, since fuzzy relational equations are special cases of the bipolar FREs, the proposed system can be also interpreted as a generalization of traditional FREs where the fuzzy compositions are generally defined by any continuous t-norm. The consistency of the continuous bipolar FREs is initially investigated and some necessary and sufficient conditions are derived for determining the feasibility of the proposed system. Subsequently, the feasible solutions set of the problem is completely characterized. It is shown that unlike FREs and those bipolar FREs defined by continuous Archimedean t-norms, the feasible solutions set of the generalized bipolar FREs is formed as the union of a finite number of compact sets that are not necessarily connected. Moreover, five techniques have also been introduced with the aim of simplifying the current problem, and then an algorithm is accordingly presented to find the feasible region of the problem. In the second step, a new class of optimization models is studied where the constraints are defined by the continuous bipolar FREs and the objective function covers a wide range of (non)linear functions such as maximum function, geometric mean function, log-sum-exp function, maximum eigenvalue of a symmetric matrix, support function of a set, etc. It is proved that the problem has a finite number of local optimal solutions and a global optimal solution can always be obtained by choosing a point having the minimum objective value compared to all the local optimal solutions. Finally, to illustrate the definitions, theorems and algorithm presented in the paper, a step-by-step example is described in several sections, whose constraints are a system of the bipolar FRES defined by Dubois-Prade t-norm (that is a continuous non-Archimedean t-norm).
Keywords: Bipolar fuzzy relational equations, continuous t-norms, Max-continuous compositions, Global optimization, Non-linear optimization.
@@ 1 Introduction
The theory of fuzzy relational equations (FRE) as a generalized version of Boolean relation equations was firstly proposed by Sanchez and was applied to problems related to the medical diagnosis [ref_43]. Pedrycz categorized and extended two ways of the generalizations of FRE in terms of sets under discussion and various operations which are taken account [39]. Since then, FRE was applied in many other fields such as fuzzy control, prediction of fuzzy systems, fuzzy decision making, fuzzy pattern recognition, image compression and reconstruction, fuzzy clustering and so on. Generally, when rules of inference are applied and their corresponding consequences are known, the problem of determining antecedents is simplified and mathematically reduced to solving an FRE [37]. Nowadays, it is well known that many of the issues associated to the body knowledge can be treated as FRE problems [38]. Because of important applications in various practical fields, many scholars have focused on the theoretical research of FRE, including its resolution approach and some specific optimization problems with FRE constraints.
@@
The solvability identification and finding set of solutions are the primary, and the most fundamental, matter concerning the FRE problems. Di Nola et al. proved that the solution set of FRE (if it is nonempty), defined by continuous max-t-norm composition is often a non-convex set. This non-convex set is completely determined by one maximum solution and a finite number of minimal solutions [8]. Such non-convexity property is one of two bottlenecks making a major contribution towards an increase in complexity of FRE-related problems, particularly, in the optimization problems subjected to a system of fuzzy relations. Another bottleneck point is concerned with detecting the minimal solutions for FREs. Chen and Wang [4] presented an algorithm for obtaining the logical representation of all minimal solutions and deduced that a polynomial-time algorithm with the ability to find all minimal solutions of FRE (with max-min composition) may not exist. Also, Markovskii showed that solving max-product FRE is closely related to the covering problem which is a type of NP-hard problem [36]. In fact, the same result holds true for a more general t-norms instead of the minimum and product operators [5, 32, 33]. Over the past decades, the solvability of FRE which is defined using different max-t compositions have been investigated by many researchers [20, 21, 23, 41, 45, 47, 48, 51, 53]. Moreover, some other researchers have worked on introducing novel concept and at times improving some of the existing theoretical aspects and applications of fuzzy relational inequalities (FRI) [18, 22, 23, 24, 29, 55]. For example, Li and Yang [29] studied FRI with addition-min composition and presented an algorithm to search for minimal solutions. They applied FRI to data transmission mechanism in a BitTorrent-like Peer-to-Peer file sharing systems. In [18], the authors focused on the study of a mixed fuzzy system formed by two fuzzy relational inequalities $A\varphi x\leq b^{1}$ and $D\varphi x\geq b^{2}$, where $\varphi$ is an operator with (closed) convex solutions.
@@
The concept of FRE was also generalized to a so-called bipolar fuzzy relation equation. Bipolarity exists widely in human understanding of information and preference [11]. Dubois and Prade provided an overview of the asymmetric bipolar representation of positive and negative information in possibility theory [12]. They showed that the possibility theory framework is convenient for handling bipolar representations, that was applied to distinguish between negative and positive information in preference modeling [11, 12]. The linear optimization of bipolar FRE was also the focus of study carried out by some researchers where FRE was defined with max-min (with application in public awareness of the products for a supplier) [15, 30], maxproduct [6] and max-Lukasiewicz composition [31, 34, 54, 57]. In [15], the concept of bipolar FRE was firstly proposed with max-min composition where the constraints are expressed as $\max_{j=1}^{n}\left\{\max\left\{\min\left\{a_{ij}^{+},x_{j}\right\},\min\left\{a_{ij}^{-},1-x_{j}\right\}\right\}\right\}$ for $i=1,\ldots,m$, where $a_{ij}^{+},a_{ij}^{-}$, $x_{j}\in[0,1]$. Similarly, in [31], the authors introduced a linear optimization problem subjected to a system of bipolar FRE defined as $X\left(A^{+},A^{-},b\right)=\left\{x\in[0,1]^{m}:x\circ A^{+}\vee\tilde{x}\circ A^{-}=b\right\}$ where $\tilde{x}_{i}=1-x_{i}$ for each component of $\tilde{x}=\left(\tilde{x}_{i}\right)_{1\times m}$ and the notations $\vee$” and $\circ$” denote max operation and the max-Lukasiewicz composition, respectively. They translated the original problem into a $0-1$ integer linear problem. In a separate, the foregoing bipolar linear optimization problem was solved by an analytical method based on the resolution and some structural properties of the feasible region (using a necessary condition for characterizing an optimal solution and a simplification process for reducing the problem) [34]. However, resolution method for obtaining the complete solution set of bipolar max-Lukasiewicz FRE was not found in the mentioned works [54]. Yang [54] studied bipolar max-Lukasiewicz FRE and showed that the complete solution set of the system is fully determined by finite conservative bipolar paths. Also, in [16], the authors studied bipolar FREs defined by max-strict t-norms. They categorized the constraints of the problem into four groups, each requiring a different approach to determine the corresponding feasible solution set.
@@
where $A^{+}=\left(a_{ij}^{+}\right)_{m\times n}$ and $A^{-}=\left(a_{ij}^{-}\right)_{m\times n}$ are fuzzy matrices and $b=\left(b_{i}\right)_{m\times 1}$ is a fuzzy vector such that $0\leq a_{ij}^{+}\leq 1$, $0\leq a_{ij}^{-}\leq 1$ and $0\leq b_{i}\leq 1$ for each $i\in\mathscr{I}=\{1,2,\ldots,m\}$ and each $j\in\mathscr{J}=\{1,2,\ldots,n\}$, respectively. Also, the sum vector 1 (a vector with each component equal to one) and the unknown vector $x=\left(x_{j}\right)_{n\times 1}$ are in $\mathbb{R}^{n}$. Moreover, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a continuous function that is non-decreasing (nonincreasing) in $x_{j}$, $\forall j\in\mathscr{J}^{+}\left(\forall j\in\mathscr{J}^{-}\right)$, where $\mathscr{J}^{+}\subseteq\mathscr{J}$, $\mathscr{J}^{-}\subseteq\mathscr{J}$, $\mathscr{J}^{+}\cap\mathscr{J}^{-}=\varnothing$ and $\mathscr{J}^{+}\cup\mathscr{J}^{-}=\mathscr{J}$. Additionally, $\varphi$ denotes an arbitrary continuous t-norm and the constraints in Problem (1) mean:
@@ 2 Preliminary definitions and properties
It should be noted that if $S\neq\varnothing$ in Definition 1, then both $l=\inf S$ and $u=\sup S$ exist and we also have the following results [18]:
Lemma 1.
@@
For the special cases of Lemma 1, we refer the reader to [16] where $\varphi$ is a continuous strict tnorm. Table A1 (see Appendix A) presents some commonly used continuous t-norms. For each t-norm in Table A1, values $l$ and $u$ discussed in Lemma 1 is presented in Table A2.
Definition 2.
@@
According to Definition 2, it is easy to verify that $S_{ij}^{+}\subseteq I_{ij}^{+}$and $S_{ij}^{-}\subseteq I_{ij}^{-},\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$. Also, the following results are directly obtained from (3) and Definition 2:
$S_{ij}^{+}$ =$=\begin{cases}{\left[l_{ij}^{+},u_{ij}^{+}\right]}&,a_{ij}^{+}\geq b_{i}\\ \varnothing&,a_{ij}^{+}<b_{i}\end{cases}\qquad$ $\quad,I_{ij}^{+}$ =$=\begin{cases}{\left[0,u_{ij}^{+}\right]}&,a_{ij}^{+}\geq b_{i}\\ {[0,1]}&,a_{ij}^{+}<b_{i}\end{cases}$ (4)
$S_{ij}^{-}$ =$=\begin{cases}{\left[1-u_{ij}^{-},1-l_{ij}^{-}\right]}&,a_{ij}^{-}\geq b_{i}\\ \varnothing&,a_{ij}^{-}<b_{i}\end{cases}$ $\quad,I_{ij}^{-}$ =$=\begin{cases}{\left[1-u_{ij}^{-},1\right]}&,a_{ij}^{-}\geq b_{i}\\ {[0,1]}&,a_{ij}^{-}<b_{i}\end{cases}$
@@
    (a) The proof is easily resulted from Definitions 2 and 3. (b) According to Definition 3, $x\in S_{ij}$ iff $\varphi\left(a_{ij}^{+},x_{j}\right)\leq b_{i}$, $\varphi\left(a_{ij}^{-},1-x_{j}\right)\leq b_{i}$ (i.e., $x\in I_{ij}^{+}\cap I_{ij}^{-}=I_{ij}$) and at least one of the two equalities $\varphi\left(a_{ij}^{+},x_{j}\right)=b_{i}$ (i.e., $x\in S_{ij}^{+}$) and $\left.\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}=b_{i}$ (i.e., $x\in S_{ij}^{-}$) holds. ∎
Based on (4) and Lemma 2, we can obtain the following results, which characterize sets $S_{ij}$ and $I_{ij}$ for all cases.
Corollary 1.
@@
Remark 1.
According to Corollary 1 and Definition 4, it is concluded that $I_{j}=\left[L_{j},U_{j}\right]$, $\forall j\in\mathscr{J}$, where
@@
In this example, =$\mathscr{I}=\{1,2,\ldots,7\}$, $\mathscr{J}=\{1,2,\ldots,9\}$ and $\varphi$ is the Dubois-Prade t-norm (that is continuous and non- Archimedean) with $\gamma=0.5$. So, according to Table A1 we have $\varphi(x,y)=T_{DP}^{0.5}(x,y)=xy/\max\{x,y,0.5\}$. For $i=1$ and $j=3$, we have $a_{13}^{+}=0.8$ and $a_{13}^{-}=b_{1}=0.7$. Hence, according to Table A2, $l_{13}^{-}=\max\left\{b_{1},\gamma\right\}=0.7$, $l_{13}^{+}=b_{1}=0.7$, $u_{13}^{-}=1$ and $u_{13}^{+}=b_{1}=0.7$. So, from Corollary 1 (d) we obtain $I_{13}=\left[1-u_{13}^{-},u_{13}^{+}\right]=[0,0.7]$ and $S_{13}=\left[1-u_{13}^{-},1-l_{13}^{-}\right]\cup\left[l_{13}^{+},u_{13}^{+}\right]=[0,0.3]\cup\{0.7\}$ (since $\left.l_{ij}^{+}=0.7>0.3=1-l_{13}^{-}\right)$. Tables 1 and 2 show all the sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. Moreover, by Definition 4 and Tables 1 and 2, it can be easily calculated that $I_{3}=\bigcap_{i\in\mathscr{I}}I_{i3}=[0.1,0.7]$ and $S_{13}^{\prime}=S_{13}\cap I_{3}=[0.1,0.3]\cup\{0.7\}$. Tables 3 and 4 show all the sets $I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively.
@@ 3 Resolution of the feasible solution set
    (a) Suppose that $x\in S\left(A^{+},A^{-},b\right)$ and $I_{j_{0}}=\varnothing$ for some $j_{0}\in\mathscr{J}$. Hence, there exists $i_{0}\in\mathscr{I}$ such that $x\notin I_{i_{0}j_{0}}$ (Definition 4). So, from Definition 3, we have $\max\left\{\varphi\left(a_{i_{0}j_{0}}^{+},x_{j_{0}}\right),\varphi\left(a_{i_{0}j_{0}}^{-},1-x_{j_{0}}\right)\right\}>b_{i_{0}}$, that implies $x\notin S_{i_{0}}$. So, from $S\left(A^{+},A^{-},b\right)=\bigcap_{i\in\mathscr{I}}S_{i}$, we obtain $x\notin S\left(A^{+},A^{-},b\right)$ that is a contradiction. (b) Assume that $x\in S\left(A^{+},A^{-},b\right)$ and there exists some $i_{0}\in\mathscr{I}$ such that $S_{i_{0}j}^{\prime}=\varnothing$, $\forall j\in\mathscr{J}$. So, $x_{j}\notin S_{i_{0}j}^{\prime}=S_{i_{0}j}\cap I_{j}$, $\forall j\in\mathscr{J}$ (Definition 4). Now, if $x_{j}\notin I_{j}$ for some $j\in\mathscr{J}$, then from Part (a) we obtain $S\left(A^{+},A^{-},b\right)\neq\varnothing$, that is a contradiction. On the other hand, if $x_{j}\notin S_{i_{0}j}^{\prime}$ and $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$, then it is concluded that $x_{j}\notin S_{i_{0}j}$, $\forall j\in\mathscr{J}$, that implies $\max\left\{\varphi\left(a_{i_{0}j}^{+},x_{j}\right),\varphi\left(a_{i_{0}j}^{-},1-x_{j}\right)\right\}<b_{i_{0}}$, $\forall j\in\mathscr{J}$ (Definition 3). Therefore, $x\notin S_{i_{0}}$, that contradicts the assumption that $x\in S\left(A^{+},A^{-},b\right)$. ∎
@@
    Suppose that $x\in[0,1]^{n}$ satisfies the conditions (I) and ( II ). So, from the condition (I) and Definitions 3 and 4, we have $x_{j}\in I_{ij}$ and therefore $\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\leq b_{i}$, $\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$ On the other hand, from the condition (II), and Definitions 3 and 4, for each $i\in\mathscr{I}$ we have $\max\left\{\varphi\left(a_{ij_{i}}^{+},x_{j_{i}}\right),\varphi\left(a_{ij_{i}}^{-},1-x_{j_{i}}\right)\right\}=b_{i}$ for some $j_{i}\in\mathscr{J}$. Consequently, for each $i\in\mathscr{I}$, it is concluded that $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\right\}=\max\left\{\varphi\left(a_{ij_{i}}^{+},x_{j_{i}}\right),\varphi\left(a_{ij_{i}}^{-},1-x_{j_{i}}\right)\right\}=b_{i}$. Thus, $x\in S_{i}$, $\forall i\in\mathscr{I}$, that implies $x\in\bigcap_{i\in\mathscr{I}}S_{i}=S\left(A^{+},A^{-},b\right)$. The converse statement is obtained by reversing the argument. ∎
Corollary 2.
@@
    The proof is resulted from Lemma 4, where $S\left(A^{+},A^{-},b\right)$ and $I_{j}$ are replaced by $S_{i}$ and $I_{ij}$, respectively. ∎
Definition 5.
@@
    $\mathscr{I}_{j}(e,i)=\{k\in\mathscr{I}:1\leq k<i$ and $e(k)=j\}$, $\forall j\in\mathscr{J}$ and $\forall i\in\mathscr{I}-\{1\}$.

### v2

@@
[…]
As a starting point, this paper develops the system of bipolar fuzzy relational equations (FRE) to the most general case, where bipolar FREs are defined by an arbitrary continuous t-norm. Due to the fact that fuzzy relational equations are special cases of bipolar FREs, the proposed system can also be viewed as a generalization of traditional FREs, in which the fuzzy composition can be defined by a continuous t-norm. In order to determine the feasibility of the proposed system, some necessary and sufficient conditions are presented for studying continuous bipolar FREs. This is followed by a complete analysis of the set of feasible solutions to the problem. Contrary to FREs and bipolar FREs defined by continuous Archimedean t-norms, the feasible solutions set of generalized bipolar FREs consists of a finite number of compact sets that are not necessarily connected. Further, five techniques have been outlined in an attempt to simplify the current problem, and then an algorithm has been presented to find the feasible region of the problem. Next, we present a class of optimization models subject to continuous bipolar FRE constraints, in which the objective function incorporates a wide range of (non)linear functions, such as maximum functions, geometric mean functions, log-sum-exp functions, maximum eigenvalues of symmetric matrices, support functions for sets, etc. Considering that the problem has a finite number of local optimal solutions, the global optimal solution can always be obtained by choosing the point with the minimum objective value among these local optimal solutions. Lastly, as a means to illustrate the definitions, theorems, and algorithms presented in the paper, a step-by-step example is presented in several sections, in which the constraints are a system of bipolar FREs defined by the Dubois-Prade t-norm, which is a continuous non-Archimedean t-norm.
[…]
Bipolar fuzzy relational equations, continuous t-norms, Max-continuous compositions, Global optimization, Non-linear optimization.
[…]
@@ 1 Introduction
The theory of fuzzy relational equations (FRE) as a generalized version of Boolean relation equations was firstly proposed by Sanchez and was applied to problems related to the medical diagnosis [42]. Pedrycz categorized and extended two ways of the generalizations of FRE in terms of sets under discussion and various operations which are taken account [37]. Since then, FRE was applied in many other fields such as fuzzy control, prediction of fuzzy systems, fuzzy decision making, fuzzy pattern recognition, image compression and reconstruction, fuzzy clustering and so on. Generally, when rules of inference are applied and their corresponding consequences are known, the problem of determining antecedents is simplified and mathematically reduced to solving an FRE [36]. Nowadays, it is well known that many of the issues associated to the body knowledge can be treated as FRE problems [38]. Because of important applications in various practical fields, many scholars have focused on the theoretical research of FRE, including its resolution approach and some specific optimization problems with FRE constraints.
@@
The solvability identification and finding set of solutions are the primary, and the most fundamental, matter concerning the FRE problems. Di Nola et al. proved that the solution set of FRE (if it is non-empty), defined by continuous max-t-norm composition is often a non-convex set. This non-convex set is completely determined by one maximum solution and a finite number of minimal solutions [8]. Such non-convexity property is one of two bottlenecks making a major contribution towards an increase in complexity of FRE-related problems, particularly, in the optimization problems subjected to a system of fuzzy relations. Another bottleneck point is concerned with detecting the minimal solutions for FREs. Chen and Wang [4] presented an algorithm for obtaining the logical representation of all minimal solutions and deduced that a polynomial-time algorithm with the ability to find all minimal solutions of FRE (with max-min composition) may not exist. Also, Markovskii showed that solving max-product FRE is closely related to the covering problem which is a type of NP-hard problem [35]. In fact, the same result holds true for a more general t-norms instead of the minimum and product operators [5, 32, 31]. Over the past decades, the solvability of FRE which is defined using different max-t compositions have been investigated by many researchers [17, 21, 15, 40, 43, 45, 46, 49, 51]. Moreover, some other researchers have worked on introducing novel concept and at times improving some of the existing theoretical aspects and applications of fuzzy relational inequalities (FRI) [20, 16, 15, 22, 27, 54]. For example, Li and Yang [27] studied FRI with addition-min composition and presented an algorithm to search for minimal solutions. They applied FRI to data transmission mechanism in a BitTorrent-like Peer-to-Peer file sharing systems. In [20], the authors focused on the study of a mixed fuzzy system formed by two fuzzy relational inequalities $A\varphi x\leq b^{1}$ and $D\varphi x\geq b^{2}$, where $\varphi$ is an operator with (closed) convex solutions.
@@
The concept of FRE was also generalized to a so-called bipolar fuzzy relation equation. Bipolarity exists widely in human understanding of information and preference [10]. Dubois and Prade provided an overview of the asymmetric bipolar representation of positive and negative information in possibility theory [11]. They showed that the possibility theory framework is convenient for handling bipolar representations, that was applied to distinguish between negative and positive information in preference modeling [10, 11]. The linear optimization of bipolar FRE was also the focus of study carried out by some researchers where FRE was defined with max-min (with application in public awareness of the products for a supplier) [14, 29], maxproduct [6] and max-Lukasiewicz composition [30, 33, 53, 55]. In [14], the concept of bipolar FRE was firstly proposed with max-min composition where the constraints are expressed as $\max_{j=1}^{n}\left\{\max\left\{\min\left\{a_{ij}^{+},x_{j}\right\},\min\left\{a_{ij}^{-},1-x_{j}\right\}\right\}\right\}$ for $i=1,\ldots,m$, where $a_{ij}^{+},a_{ij}^{-}$, $x_{j}\in[0,1]$. Similarly, in [30], the authors introduced a linear optimization problem subjected to a system of bipolar FRE defined as $X\left(A^{+},A^{-},b\right)=\left\{x\in[0,1]^{m}:x\circ A^{+}\vee\tilde{x}\circ A^{-}=b\right\}$ where $\tilde{x}_{i}=1-x_{i}$ for each component of $\tilde{x}=\left(\tilde{x}_{i}\right)_{1\times m}$ and the notations $\vee$” and $\circ$” denote max operation and the max-Lukasiewicz composition, respectively. They translated the original problem into a 010-1 integer linear problem. In a separate, the foregoing bipolar linear optimization problem was solved by an analytical method based on the resolution and some structural properties of the feasible region (using a necessary condition for characterizing an optimal solution and a simplification process for reducing the problem) [33]. However, resolution method for obtaining the complete solution set of bipolar max-Lukasiewicz FRE was not found in the mentioned works [53]. Yang [53] studied bipolar max-Lukasiewicz FRE and showed that the complete solution set of the system is fully determined by finite conservative bipolar paths. Also, in [18], the authors studied bipolar FREs defined by max-strict t-norms. They categorized the constraints of the problem into four groups, each requiring a different approach to determine the corresponding feasible solution set.
@@
where $A^{+}=\left(a_{ij}^{+}\right)_{m\times n}$ and $A^{-}=\left(a_{ij}^{-}\right)_{m\times n}$ are fuzzy matrices and $b=\left(b_{i}\right)_{m\times 1}$ is a fuzzy vector such that $0\leq a_{ij}^{+}\leq 1$, $0\leq a_{ij}^{-}\leq 1$ and $0\leq b_{i}\leq 1$ for each $i\in\mathscr{I}=\{1,2,\ldots,m\}$ and each $j\in\mathscr{J}=\{1,2,\ldots,n\}$, respectively. Also, the sum vector 1 (a vector with each component equal to one) and the unknown vector $x=\left(x_{j}\right)_{n\times 1}$ are in $\mathbb{R}^{n}$. Moreover, $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$ is a continuous function that is non-decreasing (non-increasing) in $x_{j}$, $\forall j\in\mathscr{J}^{+}\left(\forall j\in\mathscr{J}^{-}\right)$, where $\mathscr{J}^{+}\subseteq\mathscr{J}$, $\mathscr{J}^{-}\subseteq\mathscr{J}$, $\mathscr{J}^{+}\cap\mathscr{J}^{-}=\varnothing$ and $\mathscr{J}^{+}\cup\mathscr{J}^{-}=\mathscr{J}$. Additionally, $\varphi$ denotes an arbitrary continuous t-norm and the constraints in Problem (1) mean:
@@ 2 Preliminary definitions and properties
It should be noted that if $S\neq\varnothing$ in Definition 2.1, then both $l=\inf S$ and $u=\sup S$ exist and we also have the following results [20]:
Lemma 2.2.
@@
For the special cases of Lemma 2.2, we refer the reader to [18] where $\varphi$ is a continuous strict tnorm. Table A1 (see Appendix Appendix A) presents some commonly used continuous t-norms. For each t-norm in Table A1, values $l$ and $u$ discussed in Lemma 2.2 is presented in Table A2.
Definition 2.3.
@@
According to Definition 2.3, it is easy to verify that $S_{ij}^{+}\subseteq I_{ij}^{+}$and $S_{ij}^{-}\subseteq I_{ij}^{-},\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$. Also, the following results are directly obtained from (3) and Definition 2.3:
[…]
@@
(a) The proof is easily resulted from Definitions 2.3 and 2.4. (b) According to Definition 2.4, $x\in S_{ij}$ iff $\varphi\left(a_{ij}^{+},x_{j}\right)\leq b_{i}$, $\varphi\left(a_{ij}^{-},1-x_{j}\right)\leq b_{i}$ (i.e., $x\in I_{ij}^{+}\cap I_{ij}^{-}=I_{ij}$) and at least one of the two equalities $\varphi\left(a_{ij}^{+},x_{j}\right)=b_{i}$ (i.e., $x\in S_{ij}^{+}$) and $\left.\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}=b_{i}$ (i.e., $x\in S_{ij}^{-}$) holds. ∎
Based on (4) and Lemma 2.5, we can obtain the following results, which characterize sets $S_{ij}$ and $I_{ij}$ for all cases.
Corollary 2.6.
@@
Remark 2.8.
According to Corollary 2.6 and Definition 2.7, it is concluded that $I_{j}=\left[L_{j},U_{j}\right]$, $\forall j\in\mathscr{J}$, where
@@
In this example, $\mathscr{I}=\{1,2,\ldots,7\}$, $\mathscr{J}=\{1,2,\ldots,9\}$ and $\varphi$ is the Dubois-Prade t-norm (that is continuous and non- Archimedean) with $\gamma=0.5$. So, according to Table A1 we have $\varphi(x,y)=T_{DP}^{0.5}(x,y)=xy/\max\{x,y,0.5\}$. For $i=1$ and $j=3$, we have $a_{13}^{+}=0.8$ and $a_{13}^{-}=b_{1}=0.7$. Hence, according to Table A2, $l_{13}^{-}=\max\left\{b_{1},\gamma\right\}=0.7$, $l_{13}^{+}=b_{1}=0.7$, $u_{13}^{-}=1$ and $u_{13}^{+}=b_{1}=0.7$. So, from Corollary 2.6 (d) we obtain $I_{13}=\left[1-u_{13}^{-},u_{13}^{+}\right]=[0,0.7]$ and $S_{13}=\left[1-u_{13}^{-},1-l_{13}^{-}\right]\cup\left[l_{13}^{+},u_{13}^{+}\right]=[0,0.3]\cup\{0.7\}$ (since $\left.l_{ij}^{+}=0.7>0.3=1-l_{13}^{-}\right)$. Tables 1 and 2 show all the sets $I_{ij}$ and $S_{ij}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively. Moreover, by Definition 2.7 and Tables 1 and 2, it can be easily calculated that $I_{3}=\bigcap_{i\in\mathscr{I}}I_{i3}=[0.1,0.7]$ and $S_{13}^{\prime}=S_{13}\cap I_{3}=[0.1,0.3]\cup\{0.7\}$. Tables 3 and 4 show all the sets $I_{j}$ and $S_{ij}^{\prime}$ for each $i\in\mathscr{I}$ and $j\in\mathscr{J}$, respectively.
@@ 3 Resolution of the feasible solution set
(a) Suppose that $x\in S\left(A^{+},A^{-},b\right)$ and $I_{j_{0}}=\varnothing$ for some $j_{0}\in\mathscr{J}$. Hence, there exists $i_{0}\in\mathscr{I}$ such that $x\notin I_{i_{0}j_{0}}$ (Definition 2.7). So, from Definition 2.4, we have $\max\left\{\varphi\left(a_{i_{0}j_{0}}^{+},x_{j_{0}}\right),\varphi\left(a_{i_{0}j_{0}}^{-},1-x_{j_{0}}\right)\right\}>b_{i_{0}}$, that implies $x\notin S_{i_{0}}$. So, from $S\left(A^{+},A^{-},b\right)=\bigcap_{i\in\mathscr{I}}S_{i}$, we obtain $x\notin S\left(A^{+},A^{-},b\right)$ that is a contradiction. (b) Assume that $x\in S\left(A^{+},A^{-},b\right)$ and there exists some $i_{0}\in\mathscr{I}$ such that $S_{i_{0}j}^{\prime}=\varnothing$, $\forall j\in\mathscr{J}$. So, $x_{j}\notin S_{i_{0}j}^{\prime}=S_{i_{0}j}\cap I_{j}$, $\forall j\in\mathscr{J}$ (Definition 2.7). Now, if $x_{j}\notin I_{j}$ for some $j\in\mathscr{J}$, then from Part (a) we obtain $S\left(A^{+},A^{-},b\right)\neq\varnothing$, that is a contradiction. On the other hand, if $x_{j}\notin S_{i_{0}j}^{\prime}$ and $x_{j}\in I_{j}$, $\forall j\in\mathscr{J}$, then it is concluded that $x_{j}\notin S_{i_{0}j}$, $\forall j\in\mathscr{J}$, that implies $\max\left\{\varphi\left(a_{i_{0}j}^{+},x_{j}\right),\varphi\left(a_{i_{0}j}^{-},1-x_{j}\right)\right\}<b_{i_{0}}$, $\forall j\in\mathscr{J}$ (Definition 2.4). Therefore, $x\notin S_{i_{0}}$, that contradicts the assumption that $x\in S\left(A^{+},A^{-},b\right)$. ∎
@@
Suppose that $x\in[0,1]^{n}$ satisfies the conditions (I) and (II). So, from the condition (I) and Definitions 2.4 and 2.7, we have $x_{j}\in I_{ij}$ and therefore $\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\leq b_{i}$, $\forall i\in\mathscr{I}$ and $\forall j\in\mathscr{J}$ . On the other hand, from the condition (II), and Definitions 2.4 and 2.7, for each $i\in\mathscr{I}$ we have $\max\left\{\varphi\left(a_{ij_{i}}^{+},x_{j_{i}}\right),\varphi\left(a_{ij_{i}}^{-},1-x_{j_{i}}\right)\right\}=b_{i}$ for some $j_{i}\in\mathscr{J}$. Consequently, for each $i\in\mathscr{I}$, it is concluded that $\max_{j=1}^{n}\left\{\max\left\{\varphi\left(a_{ij}^{+},x_{j}\right),\varphi\left(a_{ij}^{-},1-x_{j}\right)\right\}\right\}=\max\left\{\varphi\left(a_{ij_{i}}^{+},x_{j_{i}}\right),\varphi\left(a_{ij_{i}}^{-},1-x_{j_{i}}\right)\right\}=b_{i}$. Thus, $x\in S_{i}$, $\forall i\in\mathscr{I}$, that implies $x\in\bigcap_{i\in\mathscr{I}}S_{i}=S\left(A^{+},A^{-},b\right)$. The converse statement is obtained by reversing the argument. ∎
Corollary 3.3.
@@
The proof is resulted from Lemma 3.2, where $S\left(A^{+},A^{-},b\right)$ and $I_{j}$ are replaced by $S_{i}$ and $I_{ij}$, respectively. ∎
Definition 3.4.
@@
    $\mathscr{I}_{j}(e,i)=\{k\in\mathscr{I}:1\leq k<i$ and e(k)=j}e(k)=j\}, $\forall j\in\mathscr{J}$ and $\forall i\in\mathscr{I}-\{1\}$.